OpenStudy (anonymous):

find all points on r=2sinθ where the tangent line is parallel to the ray θ=π/4

OpenStudy (dumbcow):

First to find tangent line of a function, you need the slope so we have to differentiate dr/dtheta = 2cos(theta) this has to be parallel to ray theta=pi/4 which is a straight line with slope of 1 set our slope equal to 1 2cos(theta) = 1 cos(theta) = 1/2 theta = arccos(1/2) = pi/3, 5pi/3

OpenStudy (anonymous):

Answers are : (2sinπ /8,π /8),(2sin5π /8,5π /8)

OpenStudy (anonymous):

if that helps

OpenStudy (dumbcow):

im sorry looks like i misunderstood the question what do they mean by ray of theta=pi/4 ?

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

thanks for correcting me i forgot to change back to rectangular coordinates